# Busemann Functions in the Wasserstein Space: Existence, Closed-Forms, and Applications to Slicing

### v1

@@ 1 INTRODUCTION
The Busemann function, introduced by Busemann (1955), provides a natural generalization of affine functions on non-compact metric spaces admitting geodesics which can be extended to infinity. As such, its level sets generalize the notion of affine hyperplanes and so it provides a reliable means of projecting onto geodesics. Thus, it has recently received a lot of attention in geometric Machine Learning approaches, which aim at extending classical Euclidean algorithms towards different spaces such as manifolds (Bronstein et al., 2017) for data analysis purposes.
@@
These successes in hyperbolic geometry suggest exploring the role of the Busemann function in other non-Euclidean settings. Many real-world data are best modeled as probability distributions. This is the case for instance for documents that are distributions of words (Kusner et al., 2015), single-cells (Bellazzi et al., 2021; Haviv et al., 2025b), point clouds (Haviv et al., 2025a; Geuter et al., 2025), or even images (Seguy and Cuturi, 2015). Moreover, Gaussian mixtures (Chen et al., 2018; Delon and Desolneux, 2020) or datasets with discrete labels that can be represented as mixtures of discrete distributions (Alvarez-Melis and Fusi, 2020; Bonet et al., 2025a), can be seen as datasets of probability distributions. One powerful way to endow the space of probability distributions with a metric consists in using Optimal Transport (OT) and, in particular, the Wasserstein distance (Villani, 2009). This distance allows to define the Wasserstein space, which enjoys a very rich geometry thoroughly studied in the last decades, see e.g. (Ambrosio et al., 2008; Villani, 2009; Santambrogio, 2015). Notably, it carries a formal Riemannian structure, and admits geodesics. Thus, the study of Busemann functions on the Wasserstein space is especially compelling for data analysis. A key challenge, however, is that the Wasserstein space is not geodesically complete, which prevents defining the Busemann function along every geodesic. Fortunately, for any base measure $\mu_{0}$, there is always at least one geodesic starting from $\mu_{0}$ that can be extended to infinity (Zhu et al., 2021).
@@ 2 WASSERSTEIN SPACE
In this section, we introduce the Wasserstein space and its associated Riemannian structure. Then, we study characterizations of geodesic rays on this space, i.e. geodesics that can be extended to infinity.
@@ 2.1 Wasserstein Distance
The OT problem introduces two objects of interest: the distance $\mathrm{W}_{2}$ and the optimal coupling $\gamma^{*}\in\Pi_{o}(\mu,\nu)$ solving (1). In the particular case where $\mu\in\mathcal{P}_{2,\mathrm{ac}}(\mathbb{R}^{d})$ is absolutely continuous with respect to the Lebesgue measure, it is well known by Brenier’s theorem (Brenier, 1991) that the optimal coupling $\gamma^{*}$ is unique and supported on the graph of a map $\mathrm{T}$, called the Monge map, i.e. $\gamma^{*}=(\mathrm{Id},\mathrm{T})_{\#}\mu$ with $\mathrm{T}$ satisfying $\mathrm{T}_{\#}\mu=\nu$.
@@
where $\mathcal{B}$ defines a distance between positive semi-definite matrices, known in the literature of quantum information as the Bures distance (Bhatia, 2013), and
@@
Thus, we refer to the Wasserstein distance between Gaussians as the Bures-Wasserstein distance $\mathrm{BW}$, and the space of Gaussians endowed with $\mathrm{BW}$ is called the Bures-Wasserstein space $\mathrm{BW}(\mathbb{R}^{d})$. Furthermore, the OT map between $\mu$ and $\nu$ is of the form $\mathrm{T}:x\mapsto m_{\nu}+A_{\mu}^{\nu}(x-m_{\mu})$ (Peyré et al., 2019, Remark 2.31) where
@@
If the geodesic can be extended for any $t\in\mathbb{R}$, i.e. (6) is satisfied for any $t,s\in\mathbb{R}$, it is called a geodesic line. If (6) holds for any $t,s\in\mathbb{R}_{+}$, it is called a geodesic ray (Bridson and Haefliger, 2013).
@@ 2.2 Geodesic Rays on $(\mathcal{P}_{2}(\mathbb{R}^{d}),\mathrm{W}_{2})$
Note that $\mu_{0}\in\mathcal{P}_{2,\mathrm{ac}}(\mathbb{R}^{d})$ in Proposition 1 allows leveraging Brenier’s theorem and guarantees that there exists an OT map. In the one dimensional case, we can further characterize geodesic rays starting from any $\mu_{0}\in\mathcal{P}_{2}(\mathbb{R})$ with quantile functions. Indeed, denoting $F_{0}^{-1}$ and $F_{1}^{-1}$ the quantile functions of $\mu_{0},\mu_{1}\in\mathcal{P}_{2}(\mathbb{R})$, the geodesic between $\mu_{0}$ and $\mu_{1}$ at time $t\in[0,1]$ is defined by $\mu_{t}=\big((1-t)\pi^{1}+t\pi^{2}\big)_{\#}\gamma$ with $\gamma=(F_{0}^{-1},F_{1}^{-1})_{\#}\mathrm{Unif}([0,1])$ the optimal coupling between $\mu_{0}$ and $\mu_{1}$. Then, for $F_{t}^{-1}$ the quantile of the geodesic at time $t\in[0,1]$, it is well known (see e.g. (Ambrosio et al., 2008, Equation 7.2.8)) that
@@
Let $x_{1}<\dots<x_{n}\in\mathbb{R}$, $y_{1}<\dots<y_{n}\in\mathbb{R}$, $\mu_{0}=\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}}$ and $\mu_{1}=\frac{1}{n}\sum_{i=1}^{n}\delta_{y_{i}}$. Then, the geodesic between $\mu_{0}$ and $\mu_{1}$ is a ray if and only if for all $j>i$, $y_{i}-y_{j}\leq x_{i}-x_{j}$.
@@
Note that if $\sigma_{0}^{2}=\sigma_{1}^{2}$ in Corollary 2, i.e. if the measures are translated, we recover that the geodesic is indeed a line. In the case of arbitrary 1D Gaussian distributions, we can actually obtain the largest interval over which the geodesic can be extended, see Appendix C.
@@
The conditions on $\mu_{0}$ and $\mu_{1}$ to obtain geodesic rays can be seen as guaranteeing a certain regularity along the geodesic. For instance, when $\mu_{0}\in\mathcal{P}_{2,\mathrm{ac}}(\mathbb{R}^{d})$, choosing $\mu_{1}=\nabla u_{\#}\mu_{0}$ for a 1-convex function $u$ (see Proposition 1) guarantees that the geodesic always stays in $\mathcal{P}_{2,\mathrm{ac}}(\mathbb{R}^{d})$. While it does not hold for an arbitrary $\mu_{1}\in\mathcal{P}_{2}(\mathbb{R}^{d})$, it would be possible to find the closest geodesic ray e.g. by minimizing $\mathrm{W}_{2}(\nabla u_{\#}\mu_{0},\mu_{1})$ with $u$ 1-convex (Paty et al., 2020). In the discrete case, these conditions can be interpreted in terms of particle dynamics: failure to satisfy the conditions corresponds to particle crossings. To extend arbitrary geodesics beyond $t=1$, Gallouët et al. (2025) recently proposed to solve a suitable variational problem. In particular for $\mu_{1}$ a Dirac, their solution merges particles after crossing at the Dirac.
@@ 3.1 Background on the Busemann Function
This function has attracted particular interest in geometric Machine Learning as it provides a natural generalization of hyperplanes on metric spaces. Indeed, in the particular case of Euclidean spaces, geodesic rays are of the form $\gamma(t)=t\theta$ for $\theta\in S^{d-1}=\{\theta\in\mathbb{R}^{d},\ \|\theta\|_{2}^{2}=1\}$, $t\in\mathbb{R}$, and the Busemann function is given, for any $x\in\mathbb{R}^{d}$, by $B^{\gamma}(x)=-\langle x,\theta\rangle$. Therefore, its level sets are (affine) hyperplanes. Moreover, the Busemann function provides a principled way to project a point $x\in\mathrm{X}$ on the geodesic ray $\gamma$. In fact, noticing that for any $s\in\mathbb{R}_{+}$, $B^{\gamma}\big(\gamma(s)\big)=-s$, the projection of $x\in\mathrm{X}$ on $\gamma$ is given by $P^{\gamma}(x)=\gamma\big(-B^{\gamma}(x)\big)$. In particular, all points on a level set of $B^{\gamma}$ are projected on the same point. Note however that when $B^{\gamma}(x)<0$, there is, in general, no guarantee that $\gamma\big(-B^{\gamma}(x)\big)$ belongs to the geodesic, but $B^{\gamma}(x)\in\mathbb{R}$ is always well defined, and provides a projection on $\mathbb{R}$. In a Hilbertian space (i.e. of null curvature), the Busemann projection is actually equivalent to the metric projection, i.e. $B^{\gamma}(x)=\operatorname*{argmin}_{t}\ d\big(x,\gamma(t)\big)$.
@@ 3.2 Busemann on the Wasserstein Space
First, we show that $B^{\mu}$ admits a more convenient form as an infimum over a suitable set of couplings $\Gamma(\mu_{0},\mu_{1},\nu)=\{\tilde{\gamma}\in\Pi(\mu_{0},\mu_{1},\nu),\ \pi^{1,2}_{\#}\tilde{\gamma}\in\Pi_{o}(\mu_{0},\mu_{1})\}$ between $\mu_{0},\mu_{1}$ and $\nu$, and such that the coupling between the two first marginals is optimal.
@@ 4.1 Sliced-Wasserstein Distance
for $\lambda\in\mathcal{P}(\Theta)$. In its original formulation, SW is set for $\Theta=S^{d-1}$, $\lambda=\mathcal{U}(S^{d-1})$, the uniform measure on $S^{d-1}$ and for any $\theta\in S^{d-1}$, $P^{\theta}(x)=\langle x,\theta\rangle$. However, there exist other variants with different projection schemes, e.g. convolutions for image data (Nguyen and Ho, 2022), manifold-aware projections (Bonet et al., 2023a, b, 2025b), or more general non linear projections (Kolouri et al., 2019; Chen et al., 2022).
4.2 Comparing Datasets
We consider a dataset $\mathcal{D}=\big\{(x_{i},y_{i})_{i=1}^{n}\big\}$ of pairs of samples $x_{i}\in\mathbb{R}^{d}$ associated to a label $y_{i}$ from a set of $C$ classes $\mathcal{Y}=\{1,\dots,C\}$. Class conditional distributions are given for a specific class $y$ by $\varphi(y)=\frac{1}{n_{y}}\sum_{i=1}^{n}\delta_{x_{i}}\mathbb{1}_{\{y_{i}=y\}}$. A dataset can then be represented by $\frac{1}{n}\sum_{i=1}^{n}\delta_{(x_{i},\varphi(y_{i}))}\in\mathcal{P}_{2}(\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})\big)$.
@@ 4.2 Comparing Datasets
as groundcost of the OT problem, defining the Optimal Transport Dataset Distance (OTDD), see Appendix A.4. However, OTDD is costly to compute as it requires to solve $\mathcal{O}(C^{2})$ OT problems with $n_{y}$ samples to compute the groundcost (22), and a global OT problem with $n=\sum_{y=1}^{C}n_{y}$ samples. This has led to several approximations aimed at reducing the computational cost, see e.g. (Hua et al., 2023; Liu et al., 2025).
Nguyen et al. (2025a) recently proposed a sliced distance for labeled datasets on $\mathcal{P}_{2}\big(\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})\big)$. This requires building a projection from $\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})$ to $\mathbb{R}$ to project the distribution of pairs $(x_{i},\varphi(y_{i}))$ onto a distribution in $\mathcal{P}_{2}(\mathbb{R})$. Their construction can be broken down into two projections $\mathbb{R}^{d}\to\mathbb{R}$ and $\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ using the Hierarchical Hybrid projection (Nguyen and Ho, 2024), which consists of a random linear combination with weights in the sphere. Especially, for any labeled sample $(x,y)$, their projection is of the form
@@
with $P^{\theta}:\mathbb{R}^{d}\to\mathbb{R}$, $\mathcal{M}^{\lambda}:\mathcal{P}_{2}(\mathbb{R})\to\mathbb{R}$ the moment transform projection and $\alpha\in S^{k}$. Plugging (23) into the Wasserstein distance term in (21) defines the sliced OTDD distance (SOTDD) on $\mathcal{P}_{2}\big(\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})\big)$.

### v2

@@ 1 INTRODUCTION
The Busemann function, introduced by Busemann (1955), provides a natural generalization of affine functions on non-compact metric spaces admitting geodesics which can be extended to infinity. As such, its level sets generalize the notion of affine hyperplanes and so it provides a reliable means of projecting onto geodesics. Thus, it has recently received a lot of attention in geometric Machine Learning approaches, which aim at extending classical Euclidean algorithms towards different spaces such as manifolds (Bronstein et al., 2017; Papillon et al., 2025) for data analysis purposes.
@@
These successes in hyperbolic geometry suggest exploring the role of the Busemann function in other non-Euclidean settings. Many real-world data are best modeled as probability distributions. This is the case for instance for documents that are distributions of words (Kusner et al., 2015), single-cells (Bellazzi et al., 2021; Haviv et al., 2025b), point clouds (Haviv et al., 2025a; Geuter et al., 2025), or even images (Seguy and Cuturi, 2015). Moreover, Gaussian mixtures (Chen et al., 2018; Delon and Desolneux, 2020) or datasets with discrete labels that can be represented as mixtures of discrete distributions (Alvarez-Melis and Fusi, 2020; Bonet et al., 2025a), can be seen as datasets of probability distributions. One powerful way to endow the space of probability distributions with a metric consists in using Optimal Transport (OT) and, in particular, the Wasserstein distance (Villani, 2009). This distance allows to define the Wasserstein space, which enjoys a very rich geometry thoroughly studied in the last decades, see e.g. (Ambrosio et al., 2008; Villani, 2009; Santambrogio, 2015). Notably, it carries a formal Riemannian structure, and admits geodesics. Thus, the study of Busemann functions on the Wasserstein space is especially compelling for data analysis. A key challenge, however, is that the Wasserstein space is not geodesically complete, which prevents defining the Busemann function along every geodesic. Fortunately, for any base measure $\mu_{0}$, there is always at least one geodesic starting from $\mu_{0}$ that can be extended to infinity in one direction (Zhu et al., 2021).
@@ 2 WASSERSTEIN SPACE
In this section, we introduce the Wasserstein space and its associated Riemannian structure. Then, we study characterizations of geodesic rays on this space, i.e. geodesics that can be extended to infinity in one direction. In particular, we focus on absolutely continuous probability measures, one dimensional probability measures, and Gaussians.
@@ 2.1 Wasserstein Distance
The OT problem introduces two objects of interest: the distance $\mathrm{W}_{2}$ and the optimal coupling $\gamma^{*}\in\Pi_{o}(\mu,\nu)$ solving (1). In the particular case where $\mu\in\mathcal{P}_{2,\mathrm{ac}}(\mathbb{R}^{d})$ is absolutely continuous with respect to the Lebesgue measure, it is well known by Brenier’s theorem (Brenier, 1991) that the optimal coupling $\gamma^{*}$ is unique and supported on the graph of a map $\mathrm{T}$, called the Monge map, i.e. $\gamma^{*}=(\mathrm{Id},\mathrm{T})_{\#}\mu$ with $\mathrm{T}$ satisfying $\mathrm{T}_{\#}\mu=\nu$ and $\mathrm{Id}:x\mapsto x$ the identity function.
@@
where $\mathcal{B}$ defines a distance between positive semi-definite matrices, known in the literature of quantum information as the Bures distance (Bhatia et al., 2019), and is of the form
@@
Thus, we refer to the Wasserstein distance between Gaussians as the Bures-Wasserstein distance $\mathrm{BW}$, and the space of Gaussians endowed with $\mathrm{BW}$ is called the Bures-Wasserstein space $\mathrm{BW}(\mathbb{R}^{d})$. Furthermore, the OT map between $\mu$ and $\nu$ is of the form $\mathrm{T}:x\mapsto m_{\nu}+A_{\mu}^{\nu}(x-m_{\mu})$ (Peyré and Cuturi, 2019, Remark 2.31) where
@@
If the geodesic can be extended for any $t\in\mathbb{R}$, i.e. (6) is satisfied for any $s,t\in\mathbb{R}$, it is called a geodesic line. If (6) holds for any $s,t\in\mathbb{R}_{+}$, it is called a geodesic ray (Bridson and Haefliger, 2013).
@@ 2.2 Geodesic Rays on $(\mathcal{P}_{2}(\mathbb{R}^{d}),\mathrm{W}_{2})$
Note that $\mu_{0}\in\mathcal{P}_{2,\mathrm{ac}}(\mathbb{R}^{d})$ in Proposition 1 allows leveraging Brenier’s theorem and guarantees that there exists an OT map. In the one dimensional case, we can further characterize geodesic rays starting from any $\mu_{0}\in\mathcal{P}_{2}(\mathbb{R})$ with quantile functions. Indeed, denoting $F_{0}^{-1}$ and $F_{1}^{-1}$ the quantile functions of $\mu_{0},\mu_{1}\in\mathcal{P}_{2}(\mathbb{R})$, the geodesic between $\mu_{0}$ and $\mu_{1}$ at time $t\in[0,1]$ is defined by $\mu_{t}=\big((1-t)\pi^{1}+t\pi^{2}\big)_{\#}\gamma^{*}$ with $\gamma^{*}=(F_{0}^{-1},F_{1}^{-1})_{\#}\mathrm{Unif}([0,1])$ the optimal coupling between $\mu_{0}$ and $\mu_{1}$. Then, for $F_{t}^{-1}$ the quantile of the geodesic at time $t\in[0,1]$, it is well known (see e.g. (Ambrosio et al., 2008, Equation 7.2.8)) that
@@
Let $x_{1}<\dots<x_{n}\in\mathbb{R}$, $y_{1}<\dots<y_{n}\in\mathbb{R}$, $\mu_{0}=\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}}$ and $\mu_{1}=\frac{1}{n}\sum_{i=1}^{n}\delta_{y_{i}}$. Then, the geodesic between $\mu_{0}$ and $\mu_{1}$ is a ray if and only if for all $j>i$, $y_{i}-x_{i}\leq y_{j}-x_{j}$.
Note that in the setting of Corollary 1, the geodesic is given by $\mu_{t}=\frac{1}{n}\sum_{i=1}^{n}\delta_{(1-t)x_{i}+ty_{i}}$ as points are sorted. Hence, the condition $y_{i}-x_{i}\leq y_{j}-x_{j}$ for all $j>i$ ensures that there are no crossings between particles for all time $t\geq 0$ since $x_{i}+t(y_{i}-x_{i})<x_{j}+t(y_{j}-x_{j})$. The optimal assignment between $\mu_{t}$ and $\mu_{s}$ therefore remains the same (the identity) for any $s,t\geq 0$.
@@
Note that if $\sigma_{0}^{2}=\sigma_{1}^{2}$ in Corollary 2, i.e. if the measures are translated, we recover that the geodesic is indeed a line, as it can be extended to infinity in both directions. In the case of arbitrary 1D Gaussian distributions, we can actually obtain the largest interval over which the geodesic can be extended, see Appendix C.
@@
The conditions on $\mu_{0}$ and $\mu_{1}$ to obtain geodesic rays can be seen as guaranteeing a certain regularity along the geodesic. For instance, when $\mu_{0}\in\mathcal{P}_{2,\mathrm{ac}}(\mathbb{R}^{d})$, choosing $\mu_{1}=\nabla u_{\#}\mu_{0}$ for a 1-convex function $u$ (see Proposition 1) guarantees that the geodesic always stays in $\mathcal{P}_{2,\mathrm{ac}}(\mathbb{R}^{d})$. While it does not hold for an arbitrary $\mu_{1}\in\mathcal{P}_{2}(\mathbb{R}^{d})$, it would be possible to find the closest geodesic ray e.g. by minimizing $\mathrm{W}_{2}(\nabla u_{\#}\mu_{0},\mu_{1})$ with $u$ 1-convex (Paty et al., 2020). In the discrete case, these conditions can be interpreted in terms of particle dynamics: failure to satisfy the conditions corresponds to particle crossings.
An alternative extrapolation of geodesic curves beyond $t=1$ was recently proposed by Gallouët et al. (2025) by solving a suitable variational problem. Their extrapolation coincides with the constant-speed geodesic as long as the curve remains geodesic, and therefore coincides with a true geodesic for all $t\geq 0$ when considering rays. Nonetheless, it is not the case for geodesics that are not rays. In particular for $\mu_{1}$ a Dirac, their solution merges particles after crossing at the Dirac. In this work, we only consider true geodesic rays.
We also note that the conditions to have geodesic rays are very similar to those where $\mu_{0}$ is smaller than $\mu_{1}$ in the convex order when their first moments coincide, i.e. satisfying for all $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ convex, $\int f\mathrm{d}\mu_{0}\leq\int f\mathrm{d}\mu_{1}$, see e.g. (Müller, 2001, Theorem 4 and 6) for the Gaussian case, and (Shu, 2020) for the 1D case. These conditions are also connected to projections in the convex order, see e.g. (Alfonsi and Jourdain, 2025, Proposition 3.3). We defer a more comprehensive study of these relations to future work.
@@ 3.1 Background on the Busemann Function
This function has attracted particular interest in geometric Machine Learning as it provides a natural generalization of hyperplanes on metric spaces. Indeed, in the particular case of Euclidean spaces, geodesic rays are of the form $\gamma(t)=t\theta$ for $\theta\in S^{d-1}=\{\theta\in\mathbb{R}^{d},\ \|\theta\|_{2}^{2}=1\}$, $t\in\mathbb{R}$, and the Busemann function is given, for any $x\in\mathbb{R}^{d}$, by $B^{\gamma}(x)=-\langle x,\theta\rangle$. Therefore, its level sets are (affine) hyperplanes. Moreover, the Busemann function provides a principled way to project a point $x\in\mathrm{X}$ on the geodesic ray $\gamma$. In fact, noticing that for any $s\in\mathbb{R}_{+}$, $B^{\gamma}\big(\gamma(s)\big)=-s$, the projection of $x\in\mathrm{X}$ on $\gamma$ is given by $P^{\gamma}(x)=\gamma\big(-B^{\gamma}(x)\big)$. In particular, all points on a level set of $B^{\gamma}$ are projected on the same point. Note however that when $B^{\gamma}(x)<0$, there is, in general, no guarantee that $\gamma\big(-B^{\gamma}(x)\big)$ belongs to the geodesic, but $B^{\gamma}(x)\in\mathbb{R}$ is always well defined, and provides a projection on $\mathbb{R}$. In a Hilbertian space (i.e. of null curvature), the Busemann projection is actually equivalent to the coordinate of the metric projection, i.e. $B^{\gamma}(x)=\operatorname*{argmin}_{t}\ d\big(x,\gamma(t)\big)$.
@@ 3.2 Busemann on the Wasserstein Space
First, we show that $B^{\mu}$ admits a more convenient form as an infimum over a suitable set of couplings $\Gamma(\mu_{0},\mu_{1},\nu)=\{\tilde{\gamma}\in\Pi(\mu_{0},\mu_{1},\nu),\ \pi^{1,2}_{\#}\tilde{\gamma}\in\Pi_{o}(\mu_{0},\mu_{1})\}$ between $\mu_{0},\mu_{1}$ and $\nu$, and such that the coupling between the two first marginals is optimal for (1).
@@ 4.1 Sliced-Wasserstein Distance
for $\lambda\in\mathcal{P}(\Theta)$. In its original formulation, SW is set for $\Theta=S^{d-1}$, i.e. the unit hypersphere in dimension $d$, $\lambda=\mathcal{U}(S^{d-1})$, the uniform measure on $S^{d-1}$ and for any $\theta\in S^{d-1}$, $P^{\theta}(x)=\langle x,\theta\rangle$. However, there exist other variants with different projection schemes, e.g. convolutions for image data (Nguyen and Ho, 2022), manifold-aware projections (Bonet et al., 2023a, b, 2025b), or more general non linear projections (Kolouri et al., 2019; Chen et al., 2022).
4.2 Comparing Labeled Datasets
We consider a dataset $\mathcal{D}=\big\{(x_{i},y_{i})\big\}_{i=1}^{n}$ of pairs of samples $x_{i}\in\mathbb{R}^{d}$ associated to a label $y_{i}$ from a set of $C$ classes $\mathcal{Y}=\{1,\dots,C\}$. Class conditional distributions are given for a specific class $y\in\mathcal{Y}$ by $\varphi(y)=\frac{1}{n_{y}}\sum_{i=1}^{n}\delta_{x_{i}}\mathbb{1}_{\{y_{i}=y\}}$, with $n_{y}=\sum_{i=1}^{n}\mathbb{1}_{\{y_{i}=y\}}$ the number of samples in the class $y$. A dataset can then be represented by a probability distribution over the product space $\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})$, i.e. as $\frac{1}{n}\sum_{i=1}^{n}\delta_{(x_{i},\varphi(y_{i}))}\in\mathcal{P}_{2}(\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})\big)$.
@@ 4.2 Comparing Labeled Datasets
as groundcost of an optimal transport problem on $\mathcal{P}_{2}\big(\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})\big)$, defining the Optimal Transport Dataset Distance (OTDD), see Appendix A.4. However, OTDD is costly to compute as it requires to solve $\mathcal{O}(C^{2})$ OT problems with $n_{y}$ samples to compute the groundcost (22), and a global OT problem with $n=\sum_{y=1}^{C}n_{y}$ samples. This has led to several approximations and variants aimed at reducing the computational cost, see e.g. (Hua et al., 2023; Liu et al., 2025; Nguyen et al., 2025a; Bonet et al., 2025a).
Nguyen et al. (2025a) recently proposed a sliced distance for labeled datasets on $\mathcal{P}_{2}\big(\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})\big)$. This requires building a projection from $\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})$ to $\mathbb{R}$ to project the distribution of pairs $(x_{i},\varphi(y_{i}))$ onto a distribution in $\mathcal{P}_{2}(\mathbb{R})$. Their construction can be broken down into combining two projections from $\mathbb{R}^{d}\to\mathbb{R}$ and from $\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ using the Hierarchical Hybrid projection (Nguyen and Ho, 2024), which consists of a random linear combination with weights in the sphere. Especially, for any labeled sample $(x,y)\in\mathbb{R}^{d}\times\mathcal{Y}$, their projection is of the form
@@
with $P^{\theta}:\mathbb{R}^{d}\to\mathbb{R}$ any projection on a line, $\mathcal{M}^{\lambda}:\mathcal{P}_{2}(\mathbb{R})\to\mathbb{R}$ the moment transform projection and $\alpha\in S^{k}$. Plugging (23) into the Wasserstein distance term on $\mathcal{P}_{2}(\mathbb{R})$ in (21) defines the sliced OTDD distance (SOTDD) on $\mathcal{P}_{2}\big(\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})\big)$.